Anthony invested a sum of money 6 yr ago in a savings account that has since paid interest at the rate of 8%/year compounded quarterly. His investment is now worth $61,120.62. How much did he originally invest?
Answer by lwsshak3(11628)

Anthony invested a sum of money 6 yr ago in a savings account that has since paid interest at the rate of 8%/year compounded quarterly. His investment is now worth $61,120.62. How much did he originally invest?
**
Compound interest formula: A=P(1+i)^n, i=interest rate per period, n=number of periods, A=amt after n periods.
For given problem:
i=8%/4=2%
n=4*6=24 periods(qtrs)
A=$61,120.62
P=To find
..
P=A/(1+i)^n
=A/(1+.02)^24
=61120.62/(1.02)^24
P≈38000
Anthony originally invest $38,000